OpenStudy (anonymous):

How do I find the limit of (x+3)^2 if x is approaching negative 4?

OpenStudy (anonymous):

This is a polynomial, so you can use direct substitution. Plug in -4 for x

OpenStudy (anonymous):

So, ((-4)+3)^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No foiling is required for the polynomial before plugging it in?

OpenStudy (anonymous):

you can, but it's not necessary. You'll get the same value regardless

OpenStudy (anonymous):

Okay, I just wanted to make sure I wasn't doing unnecessary steps. May I ask another quick question?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If I'm just finding the limit, I just plug the value x is approaching into the limit. So, for the limit (x/(x^2+4)) with x approaching 1. I would just plug in the value of 1 for x without any other steps?

OpenStudy (anonymous):

yes you can just plug in for that limit, but that doesn't always work, especially for rational functions

OpenStudy (anonymous):

so sometimes you will have to factor or do some other type of algebra, but it really depends on the function and the value of x it's approaching

OpenStudy (anonymous):

Do you perhaps have an example of when factoring would be required?

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2}\frac{ x+2 }{ x^2-4 }\] Plugging in -2 makes the denominator be 0. \[\lim_{x \rightarrow -2}\frac{ x+2 }{ (x+2)(x-2) }\] Cancel \[\lim_{x \rightarrow -2}\frac{ 1 }{ (x-2) }\] Now plug in to get \(-\frac{ 1 }{ 4 }\)

OpenStudy (anonymous):

I have a problem right now that wishes for me to write a simpler function that agrees with the given function at all but one point. Then to find the limit of that function. The limit as x approaches 0 of the limit (x^2+3x)/x.

OpenStudy (anonymous):

I'm not entirely sure what they mean by a simpler function, unless they would like me to factor out the x values to cancel out the x's.

OpenStudy (anonymous):

do you know what a piecewise function is?

OpenStudy (anonymous):

I haven't used a piecewise function in awhile, if you could explain it to me, I'd very much appreciate it.

OpenStudy (anonymous):

the ones that look like this, with two different functions for different parts of the domain